OpenStudy (anonymous):

Express in summation notation: 5^1+7^2+9^3+11^4+13^5+15^6+17^7

OpenStudy (anonymous):

5+49+729+1331+371293+11390625+410338673

OpenStudy (anonymous):

That's not summation notation. Thank you for giving it a shot though!

OpenStudy (anonymous):

\[\sum _{n=1}^7 (3+2 n)^n=422116015 \]From Mathematica: Table[ (3+2 n)^n, {n,7} ] = {5, 49, 729, 14641, 371293, 11390625, 410338673} Plus[ 5, 49, 729, 14641, 371293, 11390625, 410338673 ]= 422116015

OpenStudy (anonymous):

Thank you!